# Should I Trust You? Detecting Deception in Negotiations using Counterfactual RL

### v2

@@ Abstract
An increasingly prevalent socio-technical problem is people being taken in by offers that sound “too good to be true”, where persuasion and trust shape decision-making. This paper investigates how ai can help detect these deceptive scenarios. We analyze how humans strategically deceive each other in Diplomacy, a board game that requires both natural language communication and strategic reasoning. This requires extracting logical forms of proposed agreements in player communications and computing the relative rewards of the proposal using agents’ value functions. Combined with text-based features, this can improve our deception detection. Our method detects human deception with a high precision when compared to a Large Language Model approach that flags many true messages as deceptive. Future human-ai interaction tools can build on our methods for deception detection by triggering friction to give users a chance of interrogating suspicious proposals.
@@
Wichayaporn Wongkamjan1       Yanze Wang4       Feng Gu1       Denis Peskoff2 Jonathan K. Kummerfeld3       Jonathan May4       Jordan Lee Boyd-Graber1 1University of Maryland       2Princeton University       3University of Sydney 4Information Sciences Institute, University of Southern California wwongkam@umd.edu    jbg@umiacs.umd.edu
@@ 1 Friction in ai systems
Deception in natural language is a fundamental aspect of human communication, often employed as a strategic tool to mislead others through misrepresentation, omission, exaggeration, or counterfactual reasoning (Bok, 2011). From casual social interactions to high-stakes negotiations, deception influences trust, decision-making, and cooperation, making it a subject of extensive study in psychology, linguistics, and philosophy, manifesting real-world challenges such as fake news on social media  (Bade et al., 2024), misinformation  (Panda and Levitan, 2022) and adversarial communication in strategic games  (Bernard and Mickus, 2023). As artificial intelligence systems increasingly engage in human-like communication, they not only inherit but also amplify deceptive strategies, sometimes unintentionally. In ai-generated text, deception can emerge as a byproduct of optimization objectives, particularly in rlhf scenarios where agents maximize utility in multi-agent settings, sometimes at the expense of honesty Wen et al. (2025). This phenomenon has garnered significant attention across various domains, as ai deception is not confined to theoretical constructs but manifests in real-world challenges, e.g. hallucination in reasoning tasks  (Grover et al., 2024).
@@
Prior research underscores that ai-generated deceptive communication can be difficult to detect and may lead to unintended consequences when deployed in practical applications (Park et al., 2024; Sarkadi, 2024). Deceptive AI-generated text can erode trust in digital communication, amplify misinformation, and facilitate large-scale manipulation in political, financial, and social domains (Solaiman et al., 2019; Weidinger et al., 2022). Furthermore, the scalability of AI models allows deceptive content to be produced and disseminated at unprecedented rates, making manual detection impractical. To address these risks, robust mitigation strategies are necessary, including adversarial training (Perez et al., 2022), explain-ability techniques to enhance AI transparency (Danilevsky et al., 2020), and real-time detection methods leveraging linguistic and behavioral cues (Vosoughi et al., 2018).
@@
We test our detection strategies within the environment of Diplomacy, a game rich in negotiation, cooperation, and betrayal expressed through natural language. The most intriguing moments of the game arise when two players negotiate to cooperate in pursuit of their respective goals. While such agreements usually yield mutual benefits, this is not always the case—some negotiated arrangements are the result of deception, omission, or straight-up lies on the part of one player. Skilled players combat such behavior by developing the ability to recognize when an offer sounds too good to be true. Our work explores this area to raise awareness among human players when they encounter deception embedded in negotiations.
We use the value function of Cicero (Bakhtin et al., 2023), an agent trained to play Diplomacy at a human level, to detect whether a proposal is “too good to be true”. We estimate other players’ likely strategies and query Llama3 AI@Meta (2024), an llm good at general purpose semantic understanding, to determine if a message is contrary to Cicero’s expected strategy. However, Llama3 cannot precisely predict deception since it sees almost half of data as deceptive. Consequently, we decide to take a more explicit approach to modeling deception. Our contributions are as follows:
@@
1) With Theory-of-Mind-influenced deception, we identify negotiations in natural language via formal logical modeling and detect potential deceptive offers in negotiations using Cicero rl value function to generate counterfactual explanations.
@@
2) We train a BERT-based Devlin et al. (2019) classifier to predict deception using rl values and message embeddings.
@@
3) We show that our classifier is more accurate than a fine-tuned Llama3 in human lie prediction and detecting partially-deceptive negotiations.
@@ 2 Deception in the Wild
Detecting deception remains a persistent challenge, especially when it is needed for real-world problems. We can use ai, but it is hard to evaluate and requires quality feedback from humans to train models in detecting real deception. Deception in a limited space like a strategic game, e.g., Diplomacy, where nuanced persuasion and deception is required for winning, is more tractable to evaluate. A bounded example would allow us to measure the ability of an AI to improve in deception detection.
@@ 2.1 One Gains, One Loses
Deception has been studied in games that rely on trust, negotiation, and strategic misrepresentation, such as Werewolf (Chittaranjan and Hung, 2010; Hancock et al., 2017; Girlea, 2017), Poker (Lee and Hin, 2013; Palomäki et al., 2016), and Diplomacy (Niculae et al., 2015; Kramár et al., 2022). Diplomacy is a complex interplay of strategy, high-level cooperation, and subtle betrayal. The game is set on an European map, highlighting key territorial cities known as supply centers. Each of the seven players controls a country and moves units on the map, with the objective of capturing more than half of the supply centers (18 out of 34) to achieve victory. For each turn, players communicate one-to-one and then simultaneously reveal their orders for each units.
Deception plays a crucial role in gaining supply centers and, ultimately, securing a win. Cliques of players agreeing to coordinate to gain advantages over others must operate in secrecy. Deception must be undetected to be successful. If the player fails to recognize deception, they risk losing supply centers and may lose the game (Figure 1). If a player’s deception succeeds, they may gain supply centers. The challenge lies in quantifying the benefits of deception and the losses of those who are deceived. Given the significance of supply centers as a sparse scoring mechanism, we see an opportunity to integrate reinforcement learning (rl(Zinkevich et al., 2007; Brown et al., 2019) into the analysis.
rl has been extensively used to train ai agents in optimizing decision-making that maximize a reward. rl-based ai has been applied to Diplomacy Paquette et al. (2019); Anthony et al. (2020); Gray et al. (2021); Bakhtin et al. (2021), with a recent model, Cicero (Bakhtin et al., 2022), achieving competitive human-level play. This paper uses a reward model from Cicero to detect action proposals where deception is likely (Figure 2).
@@ 3.2 Deceptive Proposals
To estimate how likely a proposal is deceptive, we introduce three deceptive signs that account for different aspects of deception. First, we can measure whether a victim would get a higher reward if the proposal was not a deception (i.e., is the fake proposal from the deceiver appealing?). Second, we can measure whether a victim would get a lower reward if they believe the deception. Third, we can measure whether a deceiver would increase their future reward by deceiving the victim. These three measures called: Bait, Switch and Edge [^1] These terms come from popular culture terms around scams: a deceiver offers “bait” to attract the victim who suffers from the “switch”, leading the deceiver to profit, their “edge” in the scam. In this section, we highlight deceptive values through three hypotheses.
@@
where these can represent in a tuple as (‘A BER - KIE’,‘A MUN - RUH’, ‘F KIE - HOL’). Therefore, player $j$ can propose an action to player $i$ with multiple unit orders $\hat{a}_{i}=(\hat{a}_{i,1},\hat{a}_{i,2},\dots,\hat{a}_{i,n})$ where $n$ is a number of player $i$’s units.
We estimate how likely a proposal is deceptive by following three hypotheses when it is ‘too good to be true”. Bait, a victim perceives a greater reward if they alter a decision to follow the deceiver’s proposal and the deceiver does not actually deceive, but rather follows the plan. Assume player $i$ has a plan $a_{i}$, and player $j$ proposes $\hat{a}_{i}$ and $\hat{a}_{j}$. From the perspective of player $i$, they decide to play $\hat{a}_{i}$ because they perceive that the estimated future rewards will increase by:
@@
Switch, a victim will receive a lower reward if they follow the deceiver’s request and if the deceiver betrays the victim. Player $j$ proposes actions $\hat{a}_{i}$ and $\hat{a}_{j}$ to player $i$ where player $j$ has alternative plan $a_{j}$ to instantly stab or take advantage of player $i$. The estimated future rewards of player $i$ will decrease if player $j$ betrays player $i$. We leverage Cicero’s rl value function $u_{i}$ (Section 3.1) to formulate the first hypothesis:
@@
Edge, a deceiver will receive a better reward when a victim follows their proposal. Given the deceiver $j$’s plan $a_{j}$ and the victim’s plan $a_{i}$, if player $j$ proposes a suboptimal $\hat{a}_{i}$ to player $i$ and player $i$ falls for it. The estimated future rewards for player $j$ can increase:
@@
In short, the three hypotheses for deceptive proposals assume the victim loses, the deceiver gains and victim follows the proposal (Counterfactual RL, Figure 2). For a victim or player $i$’s plan $a_{i}$ in Equation 1 and Equation 3, we define the plan loosely as any action closest to the optimal action from player $i$’s perspective, thus sampling an action $a_{i}\sim\pi_{i}$ where $\pi_{i}$ is Cicero’s policy.
In the final step, we train a classifier using deceptive values $U_{1}$, $U_{2}$ and $U_{3}$ from Equation 1, 2 and 3. We have two main models:
1) A bert-based-uncased that takes a message and outputs text embeddings.
2) A Neural Network with three linear layers, ReLU and Sigmoid activations, that takes in text embeddings concatenated with $U_{1}$, $U_{2}$ and $U_{3}$
@@ 4 Recall-Oriented Lie Detection for Friction
This section explores deception detection and its role in creating strategic friction, i.e., deliberate decision-making in human-ai interactions. Section 4.1 analyzes human-only Diplomacy games (Peskov et al., 2020) to categorize deceptive messages and select messages for training and evaluation. Section 4.2 extends this analysis to a larger dataset within human-ai settings, testing whether our framework can introduce friction against deceptive proposals. Our goal is not to optimize $F_{1}$-Score but rather to flag possible deception for users, introducing friction to help them detect deception. In other words, our goal is to maximize recall.
@@ 4.1 Alignment to Human Lies
  • For training data, we focus on messages with human-annotated negotiations—logical form of negotiations annotated by experts—specifically for player $i$ or player $j$. In total, there are 344 messages with human-annotated negotiations containing 59 lies and 28 proposals. We sample additional 1,500 messages, though without human annotation, we retrieve logical forms of negotiations as discussed in Section 3.
  • For evaluation, we simply sample 1,000 messages containing 80 lies from the rest of dataset without any further selection.
@@ 4.2 Friction for Humans
While the data from Peskov et al. (2020) confirms our approach, CTRL-D, has a desired recall, it is small. Thus, we next test generalization on Meta’s dataset [^4] AI@Meta: Towards Human-AI Cooperation RFP of Diplomacy, which contains 40,000 games, 13 million natural language interactions from humans and Cicero players. Although these data lack thorough deception annotation, we can ex post facto validate precision through human verification.
@@ 5 Baselines for Deception
Deception is not only about the offer on the table: Niculae et al. (2015) show language changes before a betrayal occurs in Diplomacy, and Lai et al. (2020) demonstrate that this also holds true even in online reviews.
Thus, we compare our approach to language only baselines. Specifically, we implement an LLM-based baseline, using LLaMA 3.1-8B-Instruct as our primary model, to detect and mitigate deception in Diplomacy negotiations. Our objective is to determine whether a large language model can identify suspicious messages without relying on reinforcement learning (rl). We adopt two complementary criteria for evaluating deception: Direct Judgment and Alignment Judgment. Detailed prompts are provided in Figure A1.
@@
In the Direct Judgment approach, we prompt the baseline to decide whether a negotiation message is deceptive by examining its textual content together with relevant contextual information. For each negotiation round, we provide:
@@
    Current board state: A concise summary of each player’s positions and units (e.g., which territories are occupied by which units).
@@
    Negotiation message: The specific proposal, plan or statement made by player $j$ to player $i$.
@@
    Predicted Orders $\tilde{a}_{j}$: The orders Cicero would issue for player $j$ under a dialogue-free RL policy, indicating how Cicero believes player $j$ will actually move.
@@
Because Direct Judgment alone results in nearly half of the messages being flagged as deceptive, we propose Alignment Judgment, which applies a more structured, threshold-based process, to measure consistency between Cicero’s predicted orders and the orders proposed in the negotiation message.
For each player $j$, Cicero predicts orders $\tilde{a}_{j}$ using a dialogue-free RL policy. These predicted orders are then compared to the actions stated in the negotiation message. If player $j$ proposes $p_{j\to i}$ with actions $\hat{a}_{i}$ and $\hat{a}_{j}$, but Cicero predicts alternative actions $\tilde{a}_{j}$ that differ from $\hat{a}_{j}$, this inconsistency may indicate deception. The baseline is used to identify these inconsistencies. We define two sets: the number of misaligned orders $O_{m}=\sum_{k}\bm{1}[\tilde{a}_{j,k}\neq\hat{a}_{j,k}]$, and the number of aligned orders $O_{a}=\sum_{k}\bm{1}[\tilde{a}_{j,k}=\hat{a}_{j,k}]$, where $k$ is the number of the units of player $j$. If $O_{m}\geq O_{a}$, the message is classified as deceptive. Once flagged, the baseline triggers a risk alert, informing player $i$ of the misaligned orders. This prompts players to re-evaluate their decisions before finalizing their moves.

### v3

@@ Abstract
An increasingly common socio-technical problem is people being taken in by offers that sound “too good to be true”, where persuasion and trust shape decision-making. This paper investigates how ai can help detect these deceptive scenarios. We analyze how humans strategically deceive each other in Diplomacy, a board game that requires both natural language communication and strategic reasoning. This requires extracting logical forms representing proposals—agreements that players suggest during communication—and computing their relative rewards using agents’ value functions. Combined with text-based features, this can improve our deception detection. Our method detects human deception with a high precision when compared to a Large Language Model approach that flags many true messages as deceptive. Future human-ai interaction tools can build on our methods for deception detection by triggering friction to give users a chance of interrogating suspicious proposals. [^1] Code is available at CTRL-D github repository
@@
Wichayaporn Wongkamjan1       Yanze Wang4       Feng Gu1       Denis Peskoff2 Jonathan K. Kummerfeld3       Jonathan May4       Jordan Lee Boyd-Graber1 1Department of Computer Science, University of Maryland       2Northwestern University 3School of Computer Science, University of Sydney 4Information Sciences Institute, University of Southern California wwongkam@umd.edu    yanzewan@isi.edu    fgu1@umd.edu jonathan.kummerfeld@sydney.edu.au    jonmay@isi.edu    jbg@umiacs.umd.edu
@@ 1 Friction in ai systems
Deception in natural language is a fundamental aspect of human communication, often employed as a strategic tool to mislead others through misrepresentation, omission, exaggeration, or counterfactual reasoning (Bok, 2011). From casual social interactions to high-stakes negotiations, deception influences trust, decision-making, and cooperation, making it a subject of extensive study in psychology, linguistics, and philosophy, manifesting real-world challenges such as fake news on social media  (Bade et al., 2024), misinformation  (Panda and Levitan, 2022) and adversarial communication in strategic games  (Bernard and Mickus, 2023). As artificial intelligence systems increasingly engage in human-like communication, they not only inherit but also amplify deceptive strategies, sometimes unintentionally. In ai-generated text, deception can emerge as a byproduct of optimization objectives, particularly in rlhf where agents maximize utility in multi-agent settings, sometimes at the expense of honesty Wen et al. (2025). This phenomenon has garnered significant attention across various domains, as ai deception is not confined to theoretical constructs but manifests in real-world challenges, e.g. hallucination in reasoning tasks  (Grover et al., 2024).
@@
Prior research underscores that ai-generated deceptive communication can be difficult to detect and may lead to unintended consequences when deployed in practical applications (Park et al., 2024; Sarkadi, 2024). Deceptive AI-generated text can erode trust in digital communication, amplify misinformation, and facilitate large-scale manipulation in political, financial, and social domains (Solaiman et al., 2019; Weidinger et al., 2022). Furthermore, the scalability of AI models accelerates the production of deceptive content and dissemination of making manual detection impractical. To address these risks, robust mitigation strategies are necessary, including adversarial training (Perez et al., 2022), explain-ability techniques to enhance AI transparency (Danilevsky et al., 2020), and real-time detection methods leveraging linguistic and behavioral cues (Vosoughi et al., 2018).
@@
We test our detection strategies within the environment of Diplomacy, a game rich in negotiation, cooperation, and betrayal expressed through natural language. The most intriguing moments of the game arise when two players negotiate to cooperate in pursuit of their respective goals. While such agreements usually yield mutual benefits, this is not always the case—some negotiated arrangements are the result of deception, omission, or straight-up lies on the part of one player. Skilled players combat such behavior by developing the ability to recognize when an offer sounds too good to be true, whereas typical players struggle to recognize such situations Peskov et al. (2020); Wongkamjan et al. (2024); Gu et al. (2025). Our work explores this area to raise awareness among human players when they encounter deception embedded in negotiations.
We leverage the value function—an RL function that estimates how good a move and a position is—learned by Cicero (Bakhtin et al., 2023), the strongest ai agent trained to play Diplomacy at a human level, to evaluate whether a proposal is “too good to be true.” Our contributions are as follows:
  1. 1.
@@
    With Theory-of-Mind-influenced deception, we identify negotiations in natural language via formal logical modeling and detect potential deceptive offers in negotiations using the Cicero rl value function to generate counterfactual explanations.
  2. 2.
@@
    We train a BERT-based Devlin et al. (2019) classifier to predict deception using rl values and message embeddings.
  3. 3.
@@
    Our classifier is more accurate than a fine-tuned Llama3 in human lie prediction and detecting partially-deceptive negotiations.
@@ 2 Deception in the Wild
Detecting deception remains a persistent challenge, especially when it is needed for real-world problems. The lack of realistic, thorough deception training data precludes supervised ai approaches. Deception in a limited space like a strategic game, e.g., Diplomacy, where nuanced persuasion and deception is required for winning, is more tractable to evaluate. A bounded example would allow us to measure the ability of an ai to improve in deception detection.
@@ 2.1 One Gains, One Loses
Deception has been studied in games that rely on trust, negotiation, and strategic misrepresentation, such as Werewolf (Chittaranjan and Hung, 2010; Hancock et al., 2017; Girlea, 2017), Poker (Lee and Hin, 2013; Palomäki et al., 2016), and Diplomacy (Niculae et al., 2015; Kramár et al., 2022; Baldwin et al., 2025; Kulkarni et al., 2025). Diplomacy is a complex interplay of strategy, high-level cooperation, and subtle betrayal. The game is set on an European map, highlighting key territorial cities known as supply centers. Each of the seven players controls a country and moves units on the map, with the objective of capturing more than half of the supply centers (18 out of 34) to achieve victory. For each turn, players communicate one-to-one and then simultaneously reveal their orders for each units.
Deception plays a crucial role in gaining supply centers and, ultimately, securing a win. Cliques of players agreeing to coordinate to gain advantages over others must operate in secrecy. Deception must be undetected to be successful. If the player fails to recognize deception, they risk losing supply centers and may lose the game (Figure 1). If a player’s deception succeeds, they may gain supply centers. The challenge lies in quantifying the benefits of deception and the losses of those who are deceived. Given the significance of supply centers as a sparse scoring mechanism, we see an opportunity to integrate reinforcement learning (rl, Zinkevich et al., 2007; Brown et al., 2019) into the analysis.
rl has been extensively used to train ai agents in optimizing decision-making that maximize a reward. rl-based ai has been applied to Diplomacy Paquette et al. (2019); Anthony et al. (2020); Gray et al. (2021); Bakhtin et al. (2021), with a recent model, Cicero (Bakhtin et al., 2022), achieving competitive human-level play. This paper uses a value model from Cicero to estimate the expected future rewards of a proposed move, enabling the detection of proposals that are likely deceptive (Figure 2).
@@ 3.2 Deceptive Proposals
To estimate how likely a proposal is deceptive, we introduce three deceptive signs that account for different aspects of deception. First, we can measure whether a victim would get a higher reward if the proposal was not a deception (i.e., is the fake proposal from the deceiver appealing?). Second, we can measure whether a victim would get a lower reward if they believe the deception. Third, we can measure whether a deceiver would increase their future reward by deceiving the victim. We call these three measures: Bait, Switch and Edge [^2] These terms come from popular culture terms around scams: a deceiver offers “bait” to attract the victim who suffers from the “switch”, leading the deceiver to profit, their “edge” in the scam. In this section, we highlight deceptive values through three hypotheses.
@@
where these can represent in logical forms as (‘A BER - KIE’,‘A MUN - RUH’, ‘F KIE - HOL’). Player $j$ can propose an action to player $i$ with multiple unit orders, denoted $\hat{a}_{i}=(\hat{a}_{i,1},\hat{a}_{i,2},\dots,\hat{a}_{i,n})$, where $n$ is a number of player $i$’s units. We estimate how likely a proposal $p_{j\to i}$ is deceptive by following three hypotheses when it is ‘too good to be true”.
3.2.1 Bait
A victim perceives a greater reward if they alter a decision to follow the deceiver’s proposal and the deceiver does not actually deceive, but rather follows the plan. Assume player $i$ has a plan $a_{i}$, and player $j$ proposes $\hat{a}_{i}$ and $\hat{a}_{j}$. From the perspective of player $i$, they decide to play $\hat{a}_{i}$ because they perceive that the estimated future rewards will increase by:
@@ 3.2.1 Bait
3.2.2 Switch
A victim will receive a lower reward if they follow the deceiver’s request and if the deceiver betrays the victim. Player $j$ proposes actions $\hat{a}_{i}$ and $\hat{a}_{j}$ to player $i$ where player $j$ has alternative plan $a_{j}$ to instantly stab or take advantage of player $i$. The estimated future rewards of player $i$ will decrease if player $j$ betrays player $i$. We leverage Cicero’s rl value function $u_{i}$ (Section 3.1) to formulate the first hypothesis:
@@ 3.2.2 Switch
3.2.3 Edge
A deceiver will receive a better reward when a victim follows their proposal. Given the deceiver $j$’s plan $a_{j}$ and the victim’s plan $a_{i}$, if player $j$ proposes a suboptimal $\hat{a}_{i}$ to player $i$ and player $i$ falls for it. The estimated future rewards for player $j$ can increase:
@@ 3.2.3 Edge
In short, the three hypotheses for deceptive proposals assume the victim loses, the deceiver gains, and the victim follows the proposal (Counterfactual RL, Figure 2). For player $i$’s plan $a_{i}$ in Equation 1 and Equation 3, we define $a_{i}$ as the optimal action from player $i$’s perspective, thus sampling an action $a_{i}\sim\pi_{i}$ where $\pi_{i}$ is Cicero’s policy.
In the final step, we train a neural network classifier, three-layer linear layers with a Binary Cross Entropy loss by inputting two main features, text embedding and numeric features (shape=hidden_size + 3).
  • We retrieve text embedding by passing a text input to BERT, where we use a [CLS] token representation in the BERT output as a summarization of sentence (shape=hidden_size)
  • We directly use deceptive values $U_{1}$, $U_{2}$, $U_{3}$ as numeric features (shape=3).
  • We concatenate text embedding and numeric features and further pass them to the classifier and output a float value ranging from $[0,1]$ to predict deception probability.
@@ 4 Recall-Oriented Lie Detection for Friction
This section explores deception detection and its role in creating strategic friction, i.e., deliberate decision-making in human-ai interactions. Section 4.1 analyzes human-only Diplomacy games (Peskov et al., 2020) to categorize deceptive messages and select messages for training and evaluation. Section 4.2 extends this analysis to a larger dataset within human-ai settings, testing whether our framework can introduce friction against deceptive proposals. Our goal is not to optimize $F_{1}$-Score but rather to flag possible deception for users, introducing friction to help them detect deception. In other words, our goal is to examine these cases.
@@ 4.1 Alignment to Human Lies
  • For training data, we focus on messages with human-annotated negotiations—logical form of negotiations annotated by experts—specifically for player $i$ or player $j$. In total, there are 344 messages with human-annotated negotiations containing fifty-nine lies and twenty-eight proposals. We sample additional 1,500 messages, though without human annotation, we retrieve logical forms of negotiations as discussed in Section 3.
  • For evaluation, we sample 1,000 messages containing eighty lies from the rest of dataset without any further selection.
@@ 4.2 Friction for Humans
While the data from Peskov et al. (2020) confirms our approach: CTRL-D has a desired recall, it is small. Thus, we next test generalization on Meta’s data set curated from webdiplomacy.net [^5] AI@Meta: Towards Human-AI Cooperation RFP, which contains 40,000 games, 13 million natural language interactions from humans and Cicero players. Although these data lack thorough deception annotation, we can ex post facto validate precision through human verification.
@@ 5 Baselines for Deception
Deception is not only about the offer on the table: Niculae et al. (2015) show language changes before a betrayal occurs in Diplomacy, and Lai et al. (2020) demonstrate that this also holds true even in online reviews. These findings motivate the use of linguistic signals alone to detect deception, independent of game mechanics or strategic models. Thus, we compare our approach to language only baselines. Specifically, we implement an LLM-based baseline, using LLaMA 3.1-8B-Instruct (AI@Meta, 2024) as our primary model, to detect and mitigate deception in Diplomacy negotiations. Our objective is to determine whether a large language model can identify suspicious messages without relying on reinforcement learning (rl). We adopt two complementary criteria for evaluating deception: Direct Judgment and Alignment Judgment. Detailed prompts are provided in Figure A1.
@@
We prompt the baseline to decide whether a negotiation message is deceptive by examining its textual content together with relevant contextual information. For each negotiation round, we provide:
@@
    Current board state: A summary of player’s positions and units (e.g., which territories are occupied by which units).
@@
    Negotiation message: The specific proposal made by player $j$ to player $i$.
@@
    Predicted Orders $\tilde{a}_{j}$: The orders Cicero would issue for player $j$ under an RL policy, indicating how Cicero believes player $j$ will actually move.
@@
We propose Alignment Judgment which measures consistency between Cicero’s predicted orders and the orders proposed in the negotiation message. Since Cicero’s strategy model is trained to be near optimal, it is reliable and we see this as an opportunity to provide predicted orders as additional information.
For each player $j$, Cicero predicts orders $\tilde{a}_{j}$ using an RL policy. These predicted orders are compared to the actions stated in the negotiation message. If player $j$ proposes $p_{j\to i}$ with actions $\hat{a}_{i}$ and $\hat{a}_{j}$, but Cicero predicts alternative actions $\tilde{a}_{j}$ where $\tilde{a}_{j}\neq\hat{a}_{j}$, this inconsistency may indicate deception. We define two sets: the number of misaligned orders $O_{m}=\sum_{k}\bm{1}[\tilde{a}_{j,k}\neq\hat{a}_{j,k}]$, and the number of aligned orders $O_{a}=\sum_{k}\bm{1}[\tilde{a}_{j,k}=\hat{a}_{j,k}]$, where $k$ is the number of the units of player $j$. If $O_{m}\geq O_{a}$, the message is classified as deceptive and the baseline triggers a risk alert, informing player $i$ of the misaligned orders. This prompts players to re-evaluate their decisions before finalizing their moves.